# MADiff: Offline Multi-agent Learning with Diffusion Models

### v4

@@ Abstract
Diffusion model (DM) recently achieved huge success in various scenarios including offline reinforcement learning, where the diffusion planner learn to generate desired trajectories during online evaluations. However, despite the effectiveness in single-agent learning, it remains unclear how DMs can operate in multi-agent problems, where agents can hardly complete teamwork without good coordination by independently modeling each agent’s trajectories. In this paper, we propose MADiff, a novel generative multi-agent learning framework to tackle this problem. MADiff is realized with an attention-based diffusion model to model the complex coordination among behaviors of multiple agents. To the best of our knowledge, MADiff is the first diffusion-based multi-agent learning framework, which behaves as both a decentralized policy and a centralized controller. During decentralized executions, MADiff simultaneously performs teammate modeling, and the centralized controller can also be applied in multi-agent trajectory predictions. Our experiments show the superior performance of MADiff compared to baseline algorithms in a wide range of multi-agent learning tasks, which emphasizes the effectiveness of MADiff in modeling complex multi-agent interactions. Our code is available at https://github.com/zbzhu99/madiff.
**footnotetext: Equal contribution.
@@ 1 Introduction
Diffusion models (DMs) (Song and Ermon, 2019; Ho et al., 2020) have achieved remarkable success in various generative modeling tasks (Song et al., 2020; Meng et al., 2021a; Xu et al., 2022), owing to their exceptional abilities at capturing complex data distributions. Thus, some recent work is inspired to build more powerful decision models by introducing DMs into offline reinforcement learning (RL) (Janner et al., 2022; Ajay et al., 2023). Offline RL problems present significant challenges as they prohibit interactions with environments. Consequently, RL algorithms using temporal difference learning may incur severe extrapolation errors due to distributional shift between the behavior policy in the dataset and the learning policy (Fujimoto et al., 2019). Alternatively, viewing offline RL as return-conditioned trajectory generation with DMs avoid extrapolation errors and leverage the advances of DMs (Janner et al., 2022; Ajay et al., 2023).
@@
Despite its effectiveness in single-agent learning, applying the generative framework to multi-agent (MA) decision tasks remains uncertain. This is due to the need for modeling interactions and coordination among agents, while each agent makes their decisions in a decentralized manner. Consequently, merely adopting existing diffusion RL methods by treating each agent’s policies independently can result in serious inconsistencies due to a lack of proper credit assignment among agents. Current multi-agent learning (MAL) approaches typically train a centralized value function to update individual agents’ policies (Rashid et al., 2020) or use an autoregressive transformer to determine each agent’s actions (Meng et al., 2021b; Wen et al., 2022). However, without online interactions, an incorrect centralized value can lead to significant extrapolation errors (Fujimoto et al., 2019), and the transformer can only serve as an independent model for each agent.
@@
In this paper, we aim to study the potential of employing DMs to solve the above challenges in offline MAL problems. In particular, we formulate MAL as conditional generative modeling problems, where the target is to generate each agent’s trajectory, conditioned on the information of all agents (centralized case) or each agent (decentralized case). To address the aforementioned coordination challenges, we propose the first centralized-training-decentralized-execution (CTDE) diffusion framework for MA problems, named MADiff. MADiff adopts a novel attention-based DM to learn a return-conditional trajectory generation model on a reward-labeled multi-agent interaction dataset. In particular, the designed attention is computed in several latent layers of the model of each agent to fully interchange the information and integrate the global information of all agents. To model the coordination among agents, MADiff applies the attention mechanism on latent embedding for information interaction across agents. During training, MADiff performs centralized training on the joint trajectory distributions of all agents from offline datasets, including different levels of expected returns. During inference, MADiff adopts classifier-free guidance with low-temperature sampling to generate behaviors given the conditioned high expected returns, allowing for decentralized execution by predicting the behavior of other agents and generating its own behavior. Therefore, MADiff can be regarded as a principled offline MAL solution that not only serves as a decentralized policy for each agent or a centralized controller for all agents, but also includes teammate modeling without additional cost. Comprehensive experiments demonstrated superior performances of MADiff on various multi-agent learning tasks, including offline MARL and trajectory prediction.
@@ 2.3 Diffusing Decision Making
$\hat{\epsilon}:=\epsilon_{\theta}(\tau_{k},\emptyset,k)+\omega(\tau,w(y(\tau),k)-\epsilon_{\theta}(\hat{\tau}_{k},\emptyset,k))~{},$ (4)
@@ 3.2 Centralized Training Objectives
$\mathcal{L}(\theta,$ $\phi):=\sum_{i}\mathbb{E}_{(s^{i},a^{i},s^{\prime i})\in\mathcal{D}}[\|a^{i}-I_{\phi}^{i}(o^{i},o^{\prime i})\|^{2}]$ (6)
@@ 3.3 Centralized Control or Decentralized Execution
where every $\bm{\tilde{x}}_{t}\sim\mathcal{N}(\bm{0},\bm{I})$ is a noise vector sampled from the normal Gaussian. After generation, each agent obtains the action through its own inverse dynamics model following Equation (3) using the current observation $o_{t}^{i}$ and the predicted next observation $\hat{o}_{t+1}^{i}$, and takes a step in the environment. We highlight that MADiff provides an efficient way to generate joint actions and the attention module guarantees sufficient feature interactions and information interchange among agents.
Decentralized execution with teammate modeling. Compared with centralized control, a more popular and widely-adopted setting is that each agent only makes its own decision without any communication with other agents, which is what most current works (Lowe et al., 2017; Rashid et al., 2020; Pan et al., 2022) dealt with. In this case, we can only utilize the current local observation of each agent $i$ to plan its own trajectory. To this end, the initial noisy trajectory is conditioned on the current observation of the agent $i$. Similar to the centralized case, by iterative diffusion steps, we finally sample the joint state sequence based on the local observation of agent $i$ as:
@@
History-based generation. We find DMs are good at modeling the long-term joint distributions, and as a result MADiff perform better in some cases when we choose to condition on the trajectory of the past history instead of only the current observation. This implies that we replace the joint observation $\bm{o}_{t}$ in Equation (7) as the $C$-length joint history sequence $\bm{h}_{t}:=[\bm{o}_{t-C},\cdots,\bm{o}_{t-1},\bm{o}_{t}]$, and replace the independent observation $o_{t}^{i}$ in Equation (8) as the history sequence $h_{t}^{i}:=[o_{t-C}^{i},\cdots,o_{t-1}^{i},o_{t}^{i}]$ of each agent $i$. Appendix Section C illustrates how agents’ history and future trajectories are modeled by MADiff in both centralized control and decentralized execution.
@@ 4 Related Work
Multi-agent Offline RL. While offline RL has become an active research topic, only a limited number of works studied offline MARL due to the challenge of offline coordination. Jiang and Lu (2021) extended BCQ (Fujimoto et al., 2019), a single-agent offline RL algorithm with policy regularization to multi-agent; Yang et al. (2021) developed an implicit constraint approach for offline Q learning, which was found to perform particularly well in MAL tasks; Pan et al. (2022) argued the actor update tends to be trapped in local minima when the number of agents increases, and correspondingly proposed an actor regularization method named OMAR. All of these Q-learning-based methods naturally have extrapolation error problem (Fujimoto et al., 2019) in offline settings, and their solution cannot get rid of it but only mitigate some. As an alternative, MADT (Meng et al., 2021b) formulated offline MARL as return-conditioned supervised learning, and use a similar structure to a previous transformer-based offline RL work (Chen et al., 2021). However, offline MADT learns an independent model for each agent without modeling agent interactions; it relies on the gradient from centralized critics during online fine-tuning to integrate global information into each agent’s decentralized policy. MADiff not only avoids the problem of extrapolation error, but also achieves the modeling of collaborative information while allowing CTDE in a completely offline training manner.
@@ 5.1 Task Descriptions
Table 1: The average score on offline MARL tasks. Shaded columns represent our methods. The mean and standard error are computed over 3 different seeds.
@@
MPE Spread Expert 35.0 $\pm$ 2.6 104.0 $\pm$ 3.4 108.3 $\pm$ 3.3 98.2 $\pm$ 5.2 $114.9\pm 2.6$ 97.0 $\pm$ 4.9 $116.0\pm 3.5$
Md-Replay 10.0 $\pm$ 3.8 13.6 $\pm$ 5.7 15.4 $\pm$ 5.6 20.0 $\pm$ 8.4 $37.9\pm 12.3$ 29.1 $\pm$ 2.1 $43.1\pm 9.8$
Medium 31.6 $\pm$ 4.8 29.3 $\pm$ 5.5 29.3 $\pm$ 4.8 34.1 $\pm$ 7.2 47.9 $\pm$ 18.9 - $64.7\pm 9.6$ 58.0 $\pm$ 2.4
Random -0.5 $\pm$ 3.2 6.3 $\pm$ 3.5 9.8 $\pm$ 4.9 24.0 $\pm$ 9.8 $34.4\pm 5.3$ 7.2 $\pm$ 3.6 5.0 $\pm$ 2.2
Tag Expert 40.0 $\pm$ 9.6 113.0 $\pm$ 14.4 115.2 $\pm$ 12.5 93.9 $\pm$ 14.0 116.2 $\pm$ 19.8 - $123.9\pm 10.5$ $168.3\pm 19.0$
Md-Replay 0.9 $\pm$ 1.4 34.5 $\pm$ 27.8 28.7 $\pm$ 20.9 24.8 $\pm$ 17.3 47.1 $\pm$ 15.3 - $63.0\pm 6.8$ $98.8\pm 11.3$
Medium 22.5 $\pm$ 1.8 63.3 $\pm$ 20.0 65.1 $\pm$ 29.5 61.7 $\pm$ 23.1 66.7 $\pm$ 23.2 - $78.3\pm 7.6$ $133.5\pm 20.2$
Random 1.2 $\pm$ 0.8 2.2 $\pm$ 2.6 5.7 $\pm$ 3.5 5.0 $\pm$ 8.2 $11.1\pm 2.8$ 4.6 $\pm$ 2.6 10.0 $\pm$ 3.7
World Expert 33.0 $\pm$ 9.9 109.5 $\pm$ 22.8 110.3 $\pm$ 21.3 71.9 $\pm$ 28.1 110.4 $\pm$ 25.7 - $115.4\pm 11.1$ $178.9\pm 21.7$
Md-Replay 2.3 $\pm$ 1.5 12.0 $\pm$ 9.1 17.4 $\pm$ 8.1 29.6 $\pm$ 13.8 42.9 $\pm$ 19.5 - $60.3\pm 10.3$ $84.9\pm 3.9$
Medium 25.3 $\pm$ 2.0 71.9 $\pm$ 20.0 73.4 $\pm$ 9.3 58.6 $\pm$ 11.2 74.6 $\pm$ 11.5 - $124.2\pm 6.3$ $157.1\pm 6.8$
Random -2.4 $\pm$ 0.5 1.0 $\pm$ 3.2 2.8 $\pm$ 5.5 0.6 $\pm$ 2.0 $5.9\pm 5.2$ 0.7 $\pm$ 3.1 $6.1\pm 2.4$
MA Mujoco 2halfcheetah Good 6846 $\pm$ 574 - 7025 $\pm$ 439 - 1434 $\pm$ 1903 - $8254\pm 179$ $8662\pm 102$
Medium 1627 $\pm$ 187 - $2561\pm 82$ 1892 $\pm$ 220 - 2215 $\pm$ 27 $2221\pm 56$
Poor 465 $\pm$ 59 - 736 $\pm$ 72 - 384 $\pm$ 420 - $751\pm 74$ $767\pm 42$
2ant Good 2697 $\pm$ 267 - 2922 $\pm$ 194 - 464 $\pm$ 469 - $2940\pm 58$ $3105\pm 47$
Medium 1145 $\pm$ 126 - 744 $\pm$ 283 - 799 $\pm$ 186 - $1210\pm 89$ $1241\pm 30$
Poor 954 $\pm$ 80 - $1256\pm 122$ 857 $\pm$ 73 - 902 $\pm$ 24 1037 $\pm$ 24
4ant Good 2802 $\pm$ 133 - 2628 $\pm$ 971 - 344 $\pm$ 631 - $3090\pm 26$ $3087\pm 32$
Medium 1617 $\pm$ 153 - $1843\pm 494$ 929 $\pm$ 349 - 1679 $\pm$ 93 $1897\pm 44$
Poor 1033 $\pm$ 122 - 1075 $\pm$ 96 - 518 $\pm$ 112 - $1268\pm 51$ $1332\pm 45$
SMAC 3m Good 16.0 $\pm$ 1.0 18.8 $\pm$ 0.6 - $19.6\pm 0.3$ 19.0 $\pm$ 0.3 $19.6\pm 0.7$ $20.0\pm 0.0$
Medium 8.2 $\pm$ 0.8 18.1 $\pm$ 0.7 - $18.9\pm 0.7$ 15.8 $\pm$ 0.5 17.2 $\pm$ 0.7 $18.0\pm 0.7$
Poor 4.4 $\pm$ 0.1 $14.4\pm 1.2$ 5.8 $\pm$ 0.4 - 4.2 $\pm$ 0.1 8.9 $\pm$ 0.1 9.3 $\pm$ 0.6
2s3z Good 18.2 $\pm$ 0.4 $19.6\pm 0.3$ 19.0 $\pm$ 0.8 - 19.3 $\pm$ 0.3 19.4 $\pm$ 0.1 19.5 $\pm$ 0.1
Medium 12.3 $\pm$ 0.7 17.2 $\pm$ 0.6 - 14.3 $\pm$ 2.0 - 15.9 $\pm$ 0.9 $17.4\pm 0.3$ $17.7\pm 0.4$
Poor 6.7 $\pm$ 0.3 $12.1\pm 0.4$ 10.1 $\pm$ 0.7 - 7.0 $\pm$ 0.5 9.9 $\pm$ 0.2 10.8 $\pm$ 0.5
5m6m Good 16.6 $\pm$ 0.6 16.3 $\pm$ 0.9 - 13.8 $\pm$ 3.1 - 16.8 $\pm$ 0.1 $18.0\pm 1.0$ $18.2\pm 1.0$
Medium 12.4 $\pm$ 0.9 15.3 $\pm$ 0.7 - 17.0 $\pm$ 1.2 - 16.1 $\pm$ 0.2 $17.5\pm 0.6$ $18.0\pm 1.1$
Poor 7.5 $\pm$ 0.2 9.4 $\pm$ 0.4 - $10.4\pm 1.0$ 7.6 $\pm$ 0.3 8.9 $\pm$ 0.3 9.5 $\pm$ 0.7
8m Good 16.7 $\pm$ 0.4 $19.6\pm 0.3$ 11.3 $\pm$ 6.1 - 18.5 $\pm$ 0.4 19.2 $\pm$ 0.1 $20.0\pm 0.0$
Medium 10.7 $\pm$ 0.5 18.6 $\pm$ 0.5 - 16.8 $\pm$ 3.1 - 18.2 $\pm$ 0.1 $19.2\pm 0.7$ $19.5\pm 0.9$
Poor 5.3 $\pm$ 0.1 $10.8\pm 0.8$ 4.6 $\pm$ 2.4 - 4.8 $\pm$ 0.1 5.1 $\pm$ 0.1 5.2 $\pm$ 0.1

### v5

@@ Abstract
Offline reinforcement learning (RL) aims to learn policies from pre-existing datasets without further interactions, making it a challenging task. Q-learning algorithms struggle with extrapolation errors in offline settings, while supervised learning methods are constrained by model expressiveness. Recently, diffusion models (DMs) have shown promise in overcoming these limitations in single-agent learning, but their application in multi-agent scenarios remains unclear. Generating trajectories for each agent with independent DMs may impede coordination, while concatenating all agents’ information can lead to low sample efficiency. Accordingly, we propose MADiff, which is realized with an attention-based diffusion model to model the complex coordination among behaviors of multiple agents. To our knowledge, MADiff is the first diffusion-based multi-agent learning framework, functioning as both a decentralized policy and a centralized controller. During decentralized executions, MADiff simultaneously performs teammate modeling, and the centralized controller can also be applied in multi-agent trajectory predictions. Our experiments demonstrate that MADiff outperforms baseline algorithms across various multi-agent learning tasks, highlighting its effectiveness in modeling complex multi-agent interactions.
@@ 1 Introduction
Offline reinforcement learning (RL) (Fujimoto et al., 2019; Kumar et al., 2020) learns exclusively from static datasets without online interactions, enabling the effective use of pre-collected large-scale data. However, applying temporal difference (TD) learning in offline settings causes extrapolation errors (Fujimoto et al., 2019), where target value functions are evaluated on out-of-distribution actions. Sequence modeling algorithms bypass TD-learning by directly fitting the dataset distribution (Chen et al., 2021; Janner et al., 2021). Nevertheless, these methods are limited by the model’s expressiveness, making it difficult to handle diverse datasets. They also suffer from compounding errors (Xiao et al., 2019) due to autoregressive generation. Recently, diffusion models (DMs) have achieved remarkable success in various generative modeling tasks (Song and Ermon, 2019; Ho et al., 2020; Xu et al., 2022), owing to their exceptional abilities at capturing complex, high-dimensional data distributions. Their successes have also been introduced into offline RL, offering a superior modeling choice for sequence modeling algorithms (Janner et al., 2022; Ajay et al., 2023).
@@
Compared to single-agent learning, offline multi-agent learning (MAL) has been less studied and is more challenging. Since the behaviors of all agents are interrelated, each agent is required to model interactions and coordination among agents, while making decisions in a decentralized manner to achieve the goal. Current MAL approaches typically train a centralized value function to update individual agents’ policies (Rashid et al., 2020) or use an autoregressive transformer to determine each agent’s actions (Meng et al., 2021; Wen et al., 2022). However, without online interactions, an incorrect centralized value can lead to significant extrapolation errors, and the transformer can only serve as an independent model for each agent.
In this paper, we aim to study the potential of employing DMs to solve the above challenges in offline MAL problems. Merely adopting existing diffusion RL methods by using independent DMs to model each agent can result in serious inconsistencies due to a lack of proper credit assignment among agents. Another possible solution is to concatenate all agents’ information as the input and output of the DM. However, treating the agents as a single unified agent neglects the important nature of multi-agent systems. One agent may have strong correlations with only a few other agents, which makes a full feature interaction redundant. In many multi-agent systems, agents exhibit certain symmetry and can share model parameters for efficient learning (Arel et al., 2010). However, concatenating them in a fixed order breaks this symmetry, forcing the model to treat each agent differently.
@@
To address the aforementioned coordination challenges, we propose the first centralized-training-decentralized-execution (CTDE) diffusion framework for MA problems, named MADiff. MADiff adopts a novel attention-based DM to learn a return-conditional trajectory generation model on a reward-labeled multi-agent interaction dataset. In particular, the designed attention is computed in several latent layers of the model of each agent to fully interchange the information and integrate the global information of all agents. To model the coordination among agents, MADiff applies the attention mechanism on latent embedding for information interaction across agents. The attention mechanism enables the dynamic modeling of agent interactions through learned weights, while also enabling the use of a shared backbone to model each agent’s trajectory, significantly reducing the number of parameters. During training, MADiff performs centralized training on the joint trajectory distributions of all agents from offline datasets, including different levels of expected returns. During inference, MADiff adopts classifier-free guidance with low-temperature sampling to generate behaviors given the conditioned high expected returns, allowing for decentralized execution by predicting the behavior of other agents and generating its own behavior. Therefore, MADiff can be regarded as a principled offline MAL solution that not only serves as a decentralized policy for each agent or a centralized controller for all agents, but also includes teammate modeling without additional cost. Comprehensive experiments demonstrated superior performances of MADiff on various multi-agent learning tasks, including offline MARL and trajectory prediction.
@@ 2.3 Diffusing Decision Making
$\hat{\epsilon}:=\epsilon_{\theta}(\hat{\tau}_{k},\emptyset,k)+\omega(\epsilon_{\theta}(\hat{\tau}_{k},y(\tau),k)-\epsilon_{\theta}(\hat{\tau}_{k},\emptyset,k))~{},$ (4)
@@ 3.2 Centralized Training Objectives
$\mathcal{L}(\theta,$ $\phi):=\sum_{i}\mathbb{E}_{(o^{i},a^{i},o^{\prime i})\in\mathcal{D}}[\|a^{i}-I_{\phi}^{i}(o^{i},o^{\prime i})\|^{2}]$ (6)
@@ 3.3 Centralized Control or Decentralized Execution
where every $\bm{\tilde{x}}_{K,t}\sim\mathcal{N}(\bm{0},\bm{I})$ is a noise vector sampled from the normal Gaussian. After generation, each agent obtains the action through its own inverse dynamics model following Equation (3) using the current observation $o_{t}^{i}$ and the predicted next observation $\hat{o}_{t+1}^{i}$, and takes a step in the environment. We highlight that MADiff provides an efficient way to generate joint actions and the attention module guarantees sufficient feature interactions and information interchange among agents.
Decentralized execution with teammate modeling. Compared with centralized control, a more popular and widely-adopted setting is that each agent only makes its own decision without any communication with other agents, which is what most current works (Lowe et al., 2017; Rashid et al., 2020; Wang et al., 2023) dealt with. In this case, we can only utilize the current local observation of each agent $i$ to plan its own trajectory. To this end, the initial noisy trajectory is conditioned on the current observation of the agent $i$. Similar to the centralized case, by iterative diffusion steps, we finally sample the joint state sequence based on the local observation of agent $i$ as:
@@
History-based generation. We find DMs are good at modeling the long-term joint distributions, and as a result MADiff perform better in some cases when we choose to condition on the trajectory of the past history instead of only the current observation. This implies that we replace the joint observation $\bm{o}_{t}$ in Equation (7) as the $C$-length joint history sequence $\bm{h}_{t}:=[\bm{o}_{t-C},\cdots,\bm{o}_{t-1},\bm{o}_{t}]$, and replace the independent observation $o_{t}^{i}$ in Equation (8) as the history sequence $h_{t}^{i}:=[o_{t-C}^{i},\cdots,o_{t-1}^{i},o_{t}^{i}]$ of each agent $i$. Appendix Section D illustrates how agents’ history and future trajectories are modeled by MADiff in both centralized control and decentralized execution.
@@ 4 Related Work
Multi-agent Offline RL. While offline RL has become an active research topic, only a limited number of works studied offline MARL due to the challenge of offline coordination. Jiang and Lu (2021) extended BCQ (Fujimoto et al., 2019), a single-agent offline RL algorithm with policy regularization to multi-agent; Yang et al. (2021) developed an implicit constraint approach for offline Q learning, which was found to perform particularly well in MAL tasks; Pan et al. (2022) argued the actor update tends to be trapped in local minima when the number of agents increases, and correspondingly proposed an actor regularization method named OMAR. All of these Q-learning-based methods naturally have extrapolation error problem (Fujimoto et al., 2019) in offline settings, and their solution cannot get rid of it but only mitigate some. As an alternative, MADT (Meng et al., 2021) formulated offline MARL as return-conditioned supervised learning, and use a similar structure to a previous transformer-based offline RL work (Chen et al., 2021). However, offline MADT learns an independent model for each agent without modeling agent interactions; it relies on the gradient from centralized critics during online fine-tuning to integrate global information into each agent’s decentralized policy. MADiff not only avoids the problem of extrapolation error, but also achieves the modeling of collaborative information while allowing CTDE in a completely offline training manner.
Diffusion Models for Decision-Making. There is a recent line of work applying diffusion models (DMs) to decision-making problems such as RL and imitation learning. Janner et al. (2022) design a diffusion-based trajectory generation model and train a value function to sample high-rewarded trajectories. A consequent work (Ajay et al., 2023) takes conditions as inputs to the DM, thus bringing more flexibility that generates behaviors that satisfy combinations of diverse conditions. Another line of work (Wang et al., 2022; Hansen-Estruch et al., 2023; Kang et al., 2024) uses the DM as a form of policy, i.e., generating actions conditioned on states, and the training objective behaves as a regularization under the framework of TD-based offline RL algorithms. Different from the above, SynthER (Lu et al., 2024) adopts the DM to upsample the rollout data to facilitate learning of any RL algorithms. All of these existing methods focus on solving single-agent tasks. The proposed MADiff is structurally similar to Ajay et al. (2023), but includes effective modules to model agent coordination in MAL tasks.
@@ 5.1 Task Descriptions
Table 1: The average score on offline MARL tasks. Shaded columns represent our methods. The mean and standard error are computed over 5 different seeds.
@@
MPE Spread Expert 35.0 $\pm$ 2.6 104.0 $\pm$ 3.4 108.3 $\pm$ 3.3 98.2 $\pm$ 5.2 $114.9\pm 2.6$ 95.0 $\pm$ 5.3 $116.7\pm 3.0$
Md-Replay 10.0 $\pm$ 3.8 13.6 $\pm$ 5.7 15.4 $\pm$ 5.6 20.0 $\pm$ 8.4 $37.9\pm 12.3$ 30.3 $\pm$ 2.5 $42.2\pm 8.1$
Medium 31.6 $\pm$ 4.8 29.3 $\pm$ 5.5 29.3 $\pm$ 4.8 34.1 $\pm$ 7.2 47.9 $\pm$ 18.9 - $64.9\pm 7.7$ 58.2 $\pm$ 1.7
Random -0.5 $\pm$ 3.2 6.3 $\pm$ 3.5 9.8 $\pm$ 4.9 24.0 $\pm$ 9.8 $34.4\pm 5.3$ 6.9 $\pm$ 3.1 4.3 $\pm$ 2.6
Tag Expert 40.0 $\pm$ 9.6 113.0 $\pm$ 14.4 115.2 $\pm$ 12.5 93.9 $\pm$ 14.0 116.2 $\pm$ 19.8 - $120.9\pm 14.6$ $167.6\pm 18.6$
Md-Replay 0.9 $\pm$ 1.4 34.5 $\pm$ 27.8 28.7 $\pm$ 20.9 24.8 $\pm$ 17.3 47.1 $\pm$ 15.3 - $62.3\pm 9.2$ $95.0\pm 9.7$
Medium 22.5 $\pm$ 1.8 63.3 $\pm$ 20.0 65.1 $\pm$ 29.5 61.7 $\pm$ 23.1 66.7 $\pm$ 23.2 - $77.2\pm 10.4$ $132.9\pm 15.0$
Random 1.2 $\pm$ 0.8 2.2 $\pm$ 2.6 5.7 $\pm$ 3.5 5.0 $\pm$ 8.2 $11.1\pm 2.8$ 3.2 $\pm$ 4.0 10.7 $\pm$ 4.0
World Expert 33.0 $\pm$ 9.9 109.5 $\pm$ 22.8 110.3 $\pm$ 21.3 71.9 $\pm$ 28.1 110.4 $\pm$ 25.7 - $122.6\pm 14.4$ $174.0\pm 16.8$
Md-Replay 2.3 $\pm$ 1.5 12.0 $\pm$ 9.1 17.4 $\pm$ 8.1 29.6 $\pm$ 13.8 42.9 $\pm$ 19.5 - $57.1\pm 10.7$ $83.0\pm 4.4$
Medium 25.3 $\pm$ 2.0 71.9 $\pm$ 20.0 73.4 $\pm$ 9.3 58.6 $\pm$ 11.2 74.6 $\pm$ 11.5 - $123.5\pm 4.5$ $158.2\pm 6.3$
Random -2.4 $\pm$ 0.5 1.0 $\pm$ 3.2 2.8 $\pm$ 5.5 0.6 $\pm$ 2.0 $5.9\pm 5.2$ 2.0 $\pm$ 3.0 $8.1\pm 3.5$
MA Mujoco 2halfcheetah Good 6846 $\pm$ 574 - 7025 $\pm$ 439 - 1434 $\pm$ 1903 - $8246\pm 342$ $8514\pm 336$
Medium 1627 $\pm$ 187 - $2561\pm 82$ 1892 $\pm$ 220 - 2207 $\pm$ 23 $2203\pm 65$
Poor 465 $\pm$ 59 - 736 $\pm$ 72 - 384 $\pm$ 420 - $759\pm 18$ $760\pm 15$
2ant Good 2697 $\pm$ 267 - 2922 $\pm$ 194 - 464 $\pm$ 469 - $2946\pm 77$ $3069\pm 60$
Medium 1145 $\pm$ 126 - 744 $\pm$ 283 - 799 $\pm$ 186 - $1211\pm 69$ $1243\pm 37$
Poor 954 $\pm$ 80 - $1256\pm 122$ 857 $\pm$ 73 - 946 $\pm$ 66 1038 $\pm$ 26
4ant Good 2802 $\pm$ 133 - 2628 $\pm$ 971 - 344 $\pm$ 631 - $3080\pm 38$ $3068\pm 44$
Medium 1617 $\pm$ 153 - $1843\pm 494$ 929 $\pm$ 349 - 1649 $\pm$ 100 $1871\pm 52$
Poor 1033 $\pm$ 122 - 1075 $\pm$ 96 - 518 $\pm$ 112 - $1295\pm 57$ $1353\pm 44$
SMAC 3m Good 16.0 $\pm$ 1.0 18.8 $\pm$ 0.6 - $19.6\pm 0.3$ 19.1 $\pm$ 0.5 19.3 $\pm$ 0.6 $19.9\pm 0.1$
Medium 8.2 $\pm$ 0.8 18.1 $\pm$ 0.7 - $18.9\pm 0.7$ 15.8 $\pm$ 0.4 17.3 $\pm$ 0.5 $18.1\pm 0.6$
Poor 4.4 $\pm$ 0.1 $14.4\pm 1.2$ 5.8 $\pm$ 0.4 - 4.4 $\pm$ 0.3 9.6 $\pm$ 1.7 9.5 $\pm$ 0.5
2s3z Good 18.2 $\pm$ 0.4 $19.6\pm 0.3$ 19.0 $\pm$ 0.8 - 19.3 $\pm$ 0.2 $19.6\pm 0.3$ $19.7\pm 0.3$
Medium 12.3 $\pm$ 0.7 17.2 $\pm$ 0.6 - 14.3 $\pm$ 2.0 - 15.0 $\pm$ 0.6 $17.4\pm 0.2$ $17.6\pm 0.3$
Poor 6.7 $\pm$ 0.3 $12.1\pm 0.4$ 10.1 $\pm$ 0.7 - 7.0 $\pm$ 0.3 9.8 $\pm$ 0.2 10.4 $\pm$ 0.7
5m6m Good 16.6 $\pm$ 0.6 16.3 $\pm$ 0.9 - 13.8 $\pm$ 3.1 - 16.7 $\pm$ 0.1 $17.8\pm 0.8$ $18.0\pm 0.8$
Medium 12.4 $\pm$ 0.9 15.3 $\pm$ 0.7 - 17.0 $\pm$ 1.2 - 16.6 $\pm$ 0.2 $17.3\pm 0.5$ $18.0\pm 0.8$
Poor 7.5 $\pm$ 0.2 9.4 $\pm$ 0.4 - $10.4\pm 1.0$ 7.8 $\pm$ 0.4 8.9 $\pm$ 0.2 10.3 $\pm$ 1.3
8m Good 16.7 $\pm$ 0.4 $19.6\pm 0.3$ 11.3 $\pm$ 6.1 - 18.4 $\pm$ 0.3 19.2 $\pm$ 0.1 $19.8\pm 0.4$
Medium 10.7 $\pm$ 0.5 18.6 $\pm$ 0.5 - 16.8 $\pm$ 3.1 - 18.5 $\pm$ 0.3 $18.9\pm 0.9$ $19.4\pm 0.9$
Poor 5.3 $\pm$ 0.1 $10.8\pm 0.8$ 4.6 $\pm$ 2.4 - 4.7 $\pm$ 0.1 5.1 $\pm$ 0.1 5.1 $\pm$ 0.1